OpenStudy (anonymous):

Solve the system of linear equations below. 2x + 3y = 2 and x + 6y = 4

OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4 ---> x + 6y = 4 2x + 3y = 2 ----> x + 6y =4 ----> x + 6y = 4 (2-2)x + (3-2*6)y =2 -2*4 0*x + -9y = -6 -9y =-6 y=6/9 =2/3 2x + 3y = 2 2x +3 *(2/3) =2 2x +2 =2 /-2 2x=0 x=0 Solution: (0,2/3)